# Turning Sand to Gold: Recycling Data to Bridge On-Policy and Off-Policy Learning via Causal Bound

### v1

@@ Abstract
Deep reinforcement learning (DRL) agents excel in solving complex decision-making tasks across various domains. However, they often require a substantial number of training steps and a vast experience replay buffer, leading to significant computational and resource demands. To address these challenges, we introduce a novel theoretical result that leverages the Neyman-Rubin potential outcomes framework into DRL. Unlike most methods that focus on bounding the counterfactual loss, we establish a causal bound on the factual loss, which is analogous to the on-policy loss in DRL. This bound is computed by storing past value network outputs in the experience replay buffer, effectively utilizing data that is usually discarded. Extensive experiments across the Atari 2600 and MuJoCo domains on various agents, such as DQN and SAC, achieve up to 2,427% higher reward ratio, outperforming the same agents without our proposed term, and reducing the experience replay buffer size by up to 96%, significantly improving sample efficiency at negligible cost.
@@ 1 Introduction
Figure 1: Log-scaled reward improvements comparison between agents using the additional SUFT term and the baseline agents without it across 57 Atari games. The results demonstrate the superior performance of our method across the majority of games. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Left: Double DQN; Right: PPO.
@@
In DRL, we can categorize agents into two types: on-policy agents, like A3C and PPO [19, 20], and off-policy agents, such as DQN and SAC [21, 22]. On-policy agents are generally simpler and often considered first, as they provide more stable convergence by ensuring consistent behavior and target policies. At the same time, this approach requires collecting fresh samples for each optimization iteration, which may affect sample efficiency. On the other hand, off-policy agents improve sample efficiency by leveraging past experiences stored in a replay buffer. While this improves sample efficiency, it can also lead to reduced learning stability due to the mismatch between the behavior and target policies. Improving the learning stability typically requires a substantial number of training steps and a vast experience replay buffer, leading to significant computational and resource demands.
@@
To mitigate these challenges, researchers have proposed numerous strategies designed to reduce computational demands and improve learning stability. A widely adopted approach focuses on enhancing the experience replay buffer sampling, aiming to improve convergence rates by optimizing the selection of experiences [23, 24, 25, 26]. While this approach improves convergence, it still involves high computational resources and does not fully address the divergence between on-policy and off-policy methods. Other researchers have attempted to bridge the divergence by suggesting additional computable terms, e.g., [27]. Although such approaches have improved learning stability, they often require extensive data collection and involve complex techniques.
This work addresses these limitations by introducing a novel theoretical result in causal inference and incorporating it into the DRL framework through an additional off-policy evaluation (OPE) term in the agent’s standard loss function. Our approach enables us to acquire effective causal knowledge to enhance the learning process and directly address the discrepancy between on-policy and off-policy methods. Importantly, we demonstrate that reusing existing data that is usually discarded improves sample efficiency at negligible cost.
@@
We evaluate our proposed method, SUFT, on existing off-policy agents such as DQN and SAC, as well as on-policy agents like PPO, which benefit from our method due to the deviation between behavior and target policies during optimization. Throughout comprehensive experiments on the Atari 2600 and MuJoCo domains, agents using the SUFT OPE term outperform the same ones without it. Notably, we achieve up to 2,427% higher reward ratio, as shown in Figure 2, indicating an improvement in the training convergence rate compared to an identical agent without the additional term. In addition, we reduce the experience replay buffer size by up to 96% while outperforming the baseline agent that uses a 25 times larger buffer by up to 503.36%. These significant improvements highlight the potential of our method to make DRL more computationally efficient and accessible.
@@
Figure 2: Mean reward ratio comparison between agents using the additional SUFT term and the baseline agents without it across 57 Atari games and five MuJoCo environments, highlighting the profound reward gains across diverse agents and domains.
@@ 2.1 Causal Reinforcement Learning
Causality provides a valuable perspective for creating sample-efficient agents, with research focusing on three key directions: representation learning [47, 48, 49], directed exploration [50, 51], and data augmentation [52, 53]. While many works improve sample efficiency by generating new data, our method utilizes existing discarded data from past experiences, achieving sample efficiency at negligible cost. This novel strategy sets our work apart, demonstrating that DRL agents can achieve sample efficiency with marginal computational overhead.
@@ 4.2 Setup and Definitions
Here, we present the necessary mathematical definitions required to formalize our causal upper bound. Comprehensive definitions and proofs, along with a detailed reduction from the causal framework to the DRL framework, are provided in Appendix B.
To simplify the equations, we present the binary treatment case, where there is a single target treatment, denoted by $t=1$, and a single control treatment, denoted by $t=0$. In practical DRL applications, we extend the binary case to a multiple control treatment case involving a single target policy and $N$ behavior policies, reflecting the existence of experiences generated from different behavior policies in the replay buffer. The proofs and definitions for the multiple control treatment case are in Appendix B.
We employ the Neyman-Rubin potential outcomes framework and define the following: $\mathcal{X}\subset\mathbb{R}^{d}$ is the observation space. $\mathcal{Y}\subset\mathbb{R}$ is the outcome space. $\mathcal{T}=\{0,1\}$ is the binary treatment space. Under this framework, there are two potential outcomes depending on the treatment assignment: if $t=1$, the observed outcome is $y=Y_{1}$, and if $t=0$, the observed outcome is $y=Y_{0}$. We assume there exists a joint distribution $P(X,T,Y_{0},Y_{1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $P^{x}_{Y_{0}}:=P(Y_{0}|X=x)$ is the conditional distribution of potential outcomes given the control treatment. We also assume that there are no hidden confounders in the observational data, which makes the estimated treatment effect identifiable. This is formalized by assuming that the standard strong ignorability condition holds: $(Y_{0},Y_{1})\mathrel{\perp\!\!\!\perp}T|X$, and $10<p(t=1|x)<$ 1 for all $x$. Strong ignorability is a sufficient condition for the estimated treatment effect to be identifiable [70, 57]. To integrate the potential outcomes framework into the ML context, we define a neural network $\phi(x;\theta_{t}):\mathcal{X}\to\mathcal{Y}$, parameterized by weights $\theta_{t}$ corresponding to the performed treatment $t\in\mathcal{T}$.
@@
Let $\mathbf{L}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ represent a loss function. We rely on the following inequality assumption:
@@
The proof for the ${L_{1}}$ loss is in Appendix B [^1] DRL agents are often trained using $L_{2}$ loss, for which 4.1 does not apply. While the assumption is needed to prove Theorem 4.10, we report similar experimental results regardless of whether the agent was trained using $L_{1}$ or $L_{2}$..
@@
Additionally, in DRL applications, we consider the scenario of a replay buffer containing experiences generated from $N$ distinct behavior policies. The definitions for this case are in Appendix B.
@@ 4.3 Causal Bound Theorem
The proof appears in Appendix B.
The interpretation of Theorem 4.10 in the DRL framework establishes an upper bound to the on-policy loss, by the off-policy loss, the SUFT OPE term, and a constant term $\delta$. A certain difference exists in the observational data between the causal and the DRL framework, which is detailed in Appendix B.
@@
Note that since $\delta$ is independent of $Q$, it is disregarded during the gradient-based optimization of $Q$.
To the best of our knowledge, we are the first to introduce such an upper bound within the causal framework on the factual loss, adapting it to the DRL framework, thereby bounding the on-policy loss  [^2] We name this novel theoretical result the SUFT causal bound, representing the initials of Sample efficient, Upper bound, Factual loss, and Treatment effect.. Importantly, based on our causal bound theorem, the adaptation to the DRL framework is not constrained to DQN agents but can be applied to every off-policy agent with V or Q-value networks that influence its policy. This scenario is thoroughly detailed in Appendix B. Furthermore, our method can also benefit on-policy agents, such as PPO, which encounter divergence between behavior and target policies during optimization.
@@ 4.4 Recycling Value Network Outputs
To compute the SUFT OPE term in the causal bound, we need to calculate the Q-network values corresponding to both the target and the behavior policies. While the Q-values for the target policy can be easily computed using the current Q-network, the Q-network values for the old behavior policies are not addressable since storing the network weights for every behavior policy is impractical. Notably, during the agent’s training process, the Q-values matching the behavior policies are already calculated when they are used to select the action to perform. These values are calculated for the purpose of exploitation and then discarded once the action is executed. The novel contribution of our approach lies in utilizing and reusing the value network outputs in order to compute the SUFT OPE term. This allows the reuse of data that is overlooked while enhancing the loss optimization with significant causal insights in the form of the estimated treatment effect. Our approach effectively transforms overlooked data into valuable insights, metaphorically turning sand data into gold.
To obtain value network outputs in practice, we store them in the experience replay buffer. For agents employing a Q-network, only a single Q-value corresponding to the behavior policy is stored, resulting in an experience tuple of the form $\left(s,a,r,s^{\prime},Q(s,a;\theta_{\text{behavior}})\right)$, instead of $\left(s,a,r,s^{\prime}\right)$. For agents using a V-network, the state value output is stored instead, yielding an experience tuple of the form $\left(s,a,r,s^{\prime},V(s;\theta_{\text{behavior}})\right)$, instead of $\left(s,a,r,s^{\prime}\right)$.
@@ 4.5 Universal Implementation of the SUFT OPE Term
Following the formalization of the SUFT causal bound, in Section 4.3, we adopt the well-established approach of optimizing a bound on the loss, a technique commonly employed in algorithms like expectation–maximization [71] and variational inference [72]. Consequently, rather than directly optimizing the on-policy loss, we focus on optimizing the SUFT causal bound. In addition, to improve the flexibility and accuracy of the loss optimization, we incorporate a $\lambda_{TF}$ coefficient into the SUFT OPE additional term. In the case of DQN, the SUFT causal bound objective term is as follows:
@@
In value-based methods such as DQN agents, the SUFT term is seamlessly added to the standard loss function, while in Actor-Critic architectures, it is incorporated into the critic’s loss.
@@ 5 Experiments
In this section, we provide a comprehensive empirical analysis of our proposed method, emphasizing the impact of adding the SUFT OPE term to the agent’s standard loss function across a broad and diverse range of environments. In this section, we present results obtained using ${L_{2}}$ loss, allowing us to compare our method to standard DRL agents. The results for ${L_{1}}$ loss are detailed in Appendix C.
@@ 5.1 Experimental Setup
In the Atari domain, we perform a wide range of experiments using DQN [21], which serves as the foundational algorithm for numerous Q-value network agents, Vanilla DQN, and Double DQN [73], as well as PPO [20], which is a widely adopted agent that has become a standard in the field. The Vanilla DQN results are detailed in Appendix C. For the MuJoCo domain, we evaluate our method using SAC [22], a continuous control off-policy actor-critic agent known for its robust performance, and with PPO. Across these settings, we explore different values of $\lambda_{TF}$, resulting in more than 10K individual experiments. This extensive and versatile experimentation ensures statistically significant findings, providing confidence in the robustness of the results and enabling us to draw more reliable conclusions about our method’s performance. More details about tuning $\lambda_{TF}$ are in Appendix C.
@@
To evaluate SUFT’s sample efficiency, we conduct experiments with constrained environment interactions. For the off-policy agents, we use a relatively small experience replay buffer size of 4K. Each agent configuration is independently trained 10 times per environment, with the training process limited to 400K steps for the Atari domain and 1M steps for the MuJoCo domain. Performances are analyzed by taking the median smoothed training reward from the 10 trials set, ensuring a more reliable evaluation while mitigating the influence of randomness on the results. This constrained resources experimentation is designed to demonstrate our method’s sample efficiency, and it is a common technique in DRL. [74] Results for a larger buffer size are in Appendix C.
@@
Figure 3: Learning curves comparison between agents using the SUFT OPE term and the baseline agents without it across selected Atari games. The red line represents a human reward, highlighting our method’s potential to surpass humans in specific games. Top: Double DQN; Bottom: PPO.

### v2

@@ Abstract
Deep reinforcement learning (DRL) agents excel in solving complex decision-making tasks across various domains. However, they often require a substantial number of training steps and a vast experience replay buffer, leading to significant computational and resource demands. To address these challenges, we introduce a novel theoretical result that leverages the Neyman-Rubin potential outcomes framework into DRL. Unlike most methods that focus on bounding the counterfactual loss, we establish a causal bound on the factual loss, which is analogous to the on-policy loss in DRL. This bound is computed by storing past value network outputs in the experience replay buffer, effectively utilizing data that is usually discarded. Extensive experiments across the Atari 2600 and MuJoCo domains on various agents, such as DQN and SAC, achieve up to 383% higher reward ratio, outperforming the same agents without our proposed term, and reducing the experience replay buffer size by up to 96%, significantly improving sample efficiency at a negligible cost [^1] Our code is available at https://shaham-lab.github.io/TurningSandToGold/.
@@ 1 Introduction
Figure 1: Log-scaled reward improvements comparison between agents using the additional SUFT OPE term and the baseline agents without it across 57 Atari games. The results demonstrate the superior performance of our method across the majority of games. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Left: Double DQN SUFT outperforms the baseline agent in 35 out of the 40 valid games; Right: PPO SUFT outperforms the baseline agent in 39 out of the 42 valid games.
@@
In DRL, we can categorize agents into two types: on-policy agents, like A3C and PPO [19, 20], and off-policy agents, such as DQN and SAC [21, 22]. On-policy agents are generally simpler and often considered first. They interact with the environment, collect experiences using a behavior policy, and once they optimize the target policy, discard all previously collected experiences to ensure that the behavior and target policies remain identical. This consistency leads to more stable convergence but at the cost of sample inefficiency, as each optimization iteration requires collecting new data. Off-policy agents, on the other hand, allow the agent to learn from previously collected experiences, meaning it can have different behavior and target policies. This methodology improves sample efficiency by leveraging past experiences stored in a replay buffer. While this improves sample efficiency, it can also lead to reduced learning stability due to the mismatch between the behavior and target policies. As a result, improving the learning stability typically requires a substantial number of training steps and a vast experience replay buffer, leading to significant computational and resource demands.
@@
To mitigate these challenges, researchers have proposed numerous strategies designed to reduce computational demands and improve learning stability. A widely adopted approach focuses on enhancing the experience replay buffer sampling, aiming to improve convergence rates by optimizing the selection of experiences [23, 24, 25, 26]. While this approach improves convergence, it still involves high computational resources and does not fully address the divergence between on-policy and off-policy methods. Other works have attempted to bridge the divergence by suggesting additional computable terms, e.g., [27]. Although such approaches have improved learning stability, they often require extensive data collection and involve complex techniques.
This work addresses these limitations by introducing a novel theoretical result in causal inference and incorporating it into the DRL framework. Specifically, we establish a causal upper bound on the unobserved factual loss using the observed counterfactual loss and the estimated treatment effect. Within the DRL framework, we incorporate this as an upper bound for the on-policy loss using the standard off-policy loss and an off-policy evaluation (OPE) term, which quantifies the estimated value discrepancy between the behavior policy and the target policy. By incorporating this term into the optimization process, our method explicitly accounts for the causal estimated treatment effect difference between the target policy and the behavior policy. This causal perspective enhances the agent’s ability to reason about the differences between the target policy and the behavior policy used to generate the experiences, thereby mitigating the fundamental divergence between on-policy and off-policy learning. Importantly, our method reuses existing data of the behavior policy network values that are usually discarded to improve sample efficiency at a negligible cost.
@@
We evaluate our proposed method, SUFT [^2] We name this novel theoretical result the SUFT causal bound, representing the initials of Sample efficient, Upper bound, Factual loss, and Treatment effect., on existing off-policy agents such as DQN and SAC, as well as on-policy agents like PPO, which benefit from our method due to the deviation between behavior and target policies during optimization. Throughout comprehensive experiments on the Atari 2600 and MuJoCo domains, agents using the SUFT OPE term outperform the same ones without it. Notably, we achieve up to 383% higher reward ratio, as shown in Figure˜2, indicating an improvement in the training convergence rate compared to an identical agent without the additional term. In addition, we reduce the experience replay buffer size by up to 96% while outperforming the baseline agent that uses a 25 times larger buffer by up to 437.05%. These significant improvements highlight the potential of our method to make DRL more computationally efficient and accessible.
@@
Figure 2: Mean reward ratio comparison between agents using the additional SUFT OPE term and the baseline agents without it across 57 Atari games and five MuJoCo environments, highlighting the profound reward gains across diverse agents and domains.
@@ 2.1 Causal Reinforcement Learning
Causality provides a valuable perspective for creating sample-efficient agents, with research focusing on three key directions: representation learning [47, 48, 49], directed exploration [50, 51], and data augmentation [52, 53]. While many works improve sample efficiency by generating new data, our method utilizes existing discarded data from past experiences, achieving sample efficiency at a negligible cost. This novel strategy sets our work apart, demonstrating that DRL agents can achieve sample efficiency with marginal computational overhead.
@@ 4.2 Setup and Definitions
Here, we present the necessary mathematical definitions required to formalize our causal upper bound. Comprehensive definitions and proofs, along with a detailed reduction from the causal framework to the DRL framework, are provided in Appendix˜B.
To simplify the equations, we present the binary treatment case, where there is a single target treatment, denoted by $t=1$, and a single control treatment, denoted by $t=0$. In practical DRL applications, we extend the binary case to a multiple control treatment case involving a single target policy and $N$ behavior policies, reflecting the existence of experiences generated from different behavior policies in the replay buffer. The proofs and definitions for the multiple control treatment case are in Appendix˜B.
We employ the Neyman-Rubin potential outcomes framework and define the following: $\mathcal{X}\subset\mathbb{R}^{d}$ is the observation space. $\mathcal{Y}\subset\mathbb{R}$ is the outcome space. $\mathcal{T}=\{0,1\}$ is the binary treatment space. Under this framework, there are two potential outcomes depending on the treatment assignment: if $t=1$, the observed outcome is $y=Y_{1}$, and if $t=0$, the observed outcome is $y=Y_{0}$. We assume there exists a joint distribution $P(X,T,Y_{0},Y_{1})$ with the following components: $q_{t}:=\Pr(T=t)$ is the treatment probability. $P^{t}_{X}:=P(X|T=t)$ is the conditional distribution of observations. $P^{x}_{Y_{1}}:=P(Y_{1}|X=x)$ is the conditional distribution of potential outcomes given the target treatment, and $P^{x}_{Y_{0}}:=P(Y_{0}|X=x)$ is the conditional distribution of potential outcomes given the control treatment. We also assume that there are no hidden confounders in the observational data, which makes the estimated treatment effect identifiable. This is formalized by assuming that the standard strong ignorability condition holds: $(Y_{0},Y_{1})\mathrel{\perp\!\!\!\perp}T|X$, and $0<p(t=1|x)<1$ for all $x$. Strong ignorability is a sufficient condition for the estimated treatment effect to be identifiable [70, 57]. To integrate the potential outcomes framework into the ML context, we define a neural network $\phi(x;\theta_{t}):\mathcal{X}\to\mathcal{Y}$, parameterized by weights $\theta_{t}$ corresponding to the performed treatment $t\in\mathcal{T}$.
@@
$\mathcal{X}:=(\mathcal{S\times A})$ is the state-action space. Potential outcome $y\sim P^{(s,a)}_{Y_{t}}$ is the accumulated reward over $n$ time steps, starting from state $s$, taking the initial action $a$, and following policy $\pi_{t}$. The hypothesis $\phi(x;\theta_{t}):=Q(s,a;\theta_{t}):(\mathcal{S\times A})\to\mathcal{Y}$ is the Q-values network, which influences the policy $\pi_{t}$.
Let $\mathbf{L}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$ represent a loss function. We rely on the following inequality assumption, which is a variation of the inverse triangle inequality:
@@
The proof for the $L_{1}$} loss is in Appendix˜B [^3] DRL agents are often trained using $L_{2}$ loss, for which 4.1 does not hold. While the assumption is needed to prove Theorem 4.10, similar experimental results were obtained with both $L_{1}$ and $L_{2}$ losses. To provide intuition for the observed $L_{2}$} performance, we empirically examined the validity of 4.1 using $L_{2}$. This analysis is detailed in Appendix D..
@@
Additionally, in DRL applications, we consider the scenario of a replay buffer containing experiences generated from $N$ distinct behavior policies. The definitions for this case are in Appendix˜B.
@@ 4.3 Causal Bound Theorem
The proof appears in Appendix˜B.
The interpretation of Theorem˜4.10 in the DRL framework establishes an upper bound to the on-policy loss, by the off-policy loss, the SUFT OPE term, and a constant term $\delta$. A certain difference exists in the observational data between the causal and the DRL framework, which is detailed in Appendix˜B.
@@
Note that since $\delta$ is independent of $Q$, it can be omitted during the gradient-based optimization of $Q$.
To the best of our knowledge, we are the first to introduce such an upper bound within the causal framework on the factual loss, adapting it to the DRL framework, thereby bounding the on-policy loss. Importantly, based on our causal bound theorem, the adaptation to the DRL framework is not constrained to DQN agents but can be applied to every off-policy agent with V or Q-value networks that influence its policy. This scenario is thoroughly detailed in Appendix˜B. Furthermore, our method is also applicable to on-policy agents, such as PPO, which benefits from it due to the divergence between the behavior and target policies that arises during epoch optimization.
@@ 4.4 Recycling Value Network Outputs
To compute the SUFT OPE term in the causal bound, we need to calculate the Q-network values corresponding to both the target and the behavior policies. While the Q-values for the target policy can be easily computed using the current Q-network, the Q-network values for the old behavior policies are not addressable since storing the network weights for every behavior policy is impractical. Notably, during the agent’s training process, the Q-values matching the behavior policies are already calculated when they are used to select the action to perform. These values are calculated for the purpose of exploitation and then discarded once the action is executed. The novel contribution of our approach lies in utilizing and reusing the value network outputs in order to compute the SUFT OPE term. This allows the reuse of data that is overlooked while enhancing the loss optimization with significant causal insights in the form of the estimated treatment effect. Our approach effectively transforms overlooked data into valuable causal insights, metaphorically turning sand data into gold.
To obtain value network outputs in practice, we store them in the experience replay buffer. For agents employing a Q-network, only a single Q-value corresponding to the behavior policy is stored, resulting in an experience tuple of the form $\left(s,a,r,s^{\prime},Q(s,a;\theta_{\text{behavior}})\right)$, instead of the standard $\left(s,a,r,s^{\prime}\right)$. For agents using a V-network, the state value output is stored instead, yielding an experience tuple of the form $\left(s,a,r,s^{\prime},V(s;\theta_{\text{behavior}})\right)$, instead of the standard $\left(s,a,r,s^{\prime}\right)$.
@@ 4.5 Universal Implementation of the SUFT OPE Term
Following the formalization of the SUFT causal bound, in Section˜4.3, we adopt the well-established approach of optimizing a bound on the loss, a technique commonly employed in algorithms like expectation–maximization [71] and variational inference [72]. Consequently, rather than directly optimizing the on-policy loss, we focus on optimizing the SUFT causal bound. In addition, to improve the flexibility and accuracy of the loss optimization, we incorporate a $\lambda_{TF}$ coefficient into the SUFT OPE additional term. In the case of DQN, the SUFT causal bound objective term is as follows:
@@
In value-based methods such as DQN agents, the SUFT OPE term is seamlessly added to the standard loss function, while in Actor-Critic architectures, it is incorporated into the critic’s loss. We provide a detailed pseudocode of our method algorithms for both DQN and PPO in Appendix˜C.
@@ 5 Experiments
In this section, we provide a comprehensive empirical analysis of our proposed method, emphasizing the impact of adding the SUFT OPE term to the agent’s standard loss function across a broad and diverse range of environments. In this section, we present results obtained using $L_{2}$} loss, allowing us to compare our method to standard DRL agents. The results for $L_{1}$} loss are detailed in Appendix˜D.
@@ 5.1 Experimental Setup
In the Atari domain, we perform a wide range of experiments using DQN [21], which serves as the foundational algorithm for numerous Q-value network agents, Vanilla DQN, and Double DQN [73], as well as PPO [20], which is a widely adopted agent that has become a standard in the field. The Vanilla DQN results are detailed in Appendix˜D. For the MuJoCo domain, we evaluate our method using SAC [22], a continuous control off-policy actor-critic agent known for its robust performance, and with PPO. Across these settings, we explore different values of $\lambda_{TF}$, resulting in more than 10K individual experiments. This extensive and versatile experimentation ensures statistically significant findings, providing confidence in the robustness of the results and enabling us to draw more reliable conclusions about our method’s performance. More details about tuning $\lambda_{TF}$ are in Appendix˜D.
@@
To evaluate SUFT’s sample efficiency, we conduct experiments with constrained environment interactions. For the off-policy agents, we use a relatively small experience replay buffer size of 4K. Each agent configuration is independently trained 10 times per environment, with the training process limited to 400K steps in the Atari domain and 1M steps in the MuJoCo domain. Performances are analyzed by taking the median smoothed training reward from the 10-trial set, ensuring a more reliable evaluation while mitigating the influence of randomness on the results. This constrained resources experimentation is designed to demonstrate our method’s sample efficiency, and it is a common technique in DRL. [74] Results for a larger buffer size are in Appendix˜D.
@@
Figure 3: Learning curves comparison between agents using the SUFT OPE term and the baseline agents without it across selected Atari games. The red line indicates human-level performance, showing that SUFT not only surpasses the baseline agent but can even exceed human rewards. Top: Double DQN; Bottom: PPO.